# On the Utility of Equivariance and Symmetry Breaking in Deep Learning Architectures on Point Clouds

### v1

@@ 4 Experiments
6 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SE}(3)$ $85.52_{\pm 0.13}$ $85.79_{\pm 0.10}$ $64.25_{\pm 0.30}$ $\sim 5\;\;$
7 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SE}(3)$ $84.35_{\pm 0.12}$ $84.74_{\pm 0.05}$ $\sim 5\;\;$
➟  8 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SE}(3)$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.69}_{\pm 0.10}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.81}_{\pm 0.06}$ $\sim 5\;\;$
9 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $84.17_{\pm 0.12}$ $84.55_{\pm 0.17}$ $\sim 5\;\;$
10 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $85.28_{\pm 0.09}$ $85.61_{\pm 0.03}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}62.09}_{\pm 1.04}$ $\sim 5\;\;$
11 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $84.44_{\pm 0.16}$ $84.68_{\pm 0.07}$ $\sim 5\;\;$
12 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $85.48_{\pm 0.07}$ $85.80_{\pm 0.08}$ $\sim 5\;\;$
@@
14 $\mathbb{R}^{3}\times S^{2}$ none $85.82_{\pm 0.11}$ $38.97_{\pm 0.83}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}63.69}_{\pm 1.17}$ $\sim 5$
15 $\mathbb{R}^{3}\times S^{2}$ none $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.69}_{\pm 0.18}$ $36.08_{\pm 0.75}$ $\sim 5\;\;$
@@
16 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}85.26_{\pm 0.01}}85.38_{\pm 0.05}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}31.41_{\pm 0.86}}32.78_{\pm 0.59}\end{matrix}$ $67.49_{\pm 0.41}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 10\end{matrix}$
17 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.82}_{\pm 0.10}}85.71_{\pm 0.17}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}35.42}_{\pm 0.98}}32.25_{\pm 0.07}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 10\end{matrix}$
18 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}85.51_{\pm 0.06}}85.52_{\pm 0.09}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}32.79_{\pm 0.74}}31.11_{\pm 0.15}\end{matrix}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}65.06}_{\pm 1.05}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 10\end{matrix}$
19 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}85.66_{\pm 0.02}}85.16_{\pm 0.06}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}33.55_{\pm 0.49}}30.62_{\pm 0.21}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 10\end{matrix}$
Reference methods from literature
@@
GeomGCNN [^3] GeomGCNN is trained with 1024 points instead of 2048. - - - - - - - 89.10 - - -
4.2 Molecular property prediction and Molecule generation (discovery)
For predicting molecular properties and generating molecules, we use QM9 (Ramakrishnan R., 2014), a dataset which consists of 130k small molecules and their 3-dimensional coordinates, along with molecular properties, integer-valued atom charges, and atom coordinates. It contains up to 9 heavy atoms and 29 atoms including hydrogens. We use the train/val/test partitions introduced in Gilmer et al. (2017), which consists of 100K/18K/13K samples respectively for each partition. We evaluate the prediction of molecular properties using the MAE metric and compare these with EGNN (Satorras et al., 2021), Dimenet++ (Gasteiger et al., 2022) and SE(3)- Transformer Fuchs et al. (2020)
[…]
Table 2: Ablation study on QM9 for property prediction task using mean absolute error (MAE) as the performance metric for three properties mu, alpha and $\epsilon_{HOMO}$, as well as molecule generation task (discovery) using atom stability, molecule stability and discover as performance metric.
@@ 4.2 Molecular property prediction and Molecule generation (discovery)
Model Variation Type ! Coordinates as Scalars ! Coordinates as Vectors Effective Equivariance MAE $\mu$ (D) MAE $\alpha$ ($a_{o}^{3}$) MAE $\epsilon_{HOMO}$ (meV) Stab Atom % Stab Mol % Discover % Normalized Epoch Time
1 $\mathbb{R}^{3}$ $\mathrm{SE}(3)$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0185}_{\pm 0.0005}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}0.0609_{\pm 0.0001}}0.1196\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}24.21_{\pm 0.20}}31.87\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 20\end{matrix}$
2 $\mathbb{R}^{3}$ none $0.0183_{\pm 0.0003}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0544}_{\pm 0.0001}}0.1195\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}23.17}_{\pm 0.3}}32.21\end{matrix}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}97.50}_{\pm 0.68}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}77.01}_{\pm 1.14}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}90.18}_{\pm 0.87}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 20\end{matrix}$
@@
➟  4 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $0.0107_{\pm 0.001}$ $0.0397_{\pm 0.0011}$ $18.80_{\pm 3.24}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}99.29}_{\pm.02}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}92.34}_{\pm.20}$ $91.13_{\pm.35}$ $\sim 10$
5 $\mathbb{R}^{3}\times S^{2}$ none $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009}_{\pm 0.0007}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038}_{\pm 0.005}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}18.42}_{\pm 0.31}$ $97.77_{\pm.02}$ $91.11_{\pm.53}$ $90.42_{\pm.84}$ $\sim 10$
@@
6 $\mathbb{R}^{3}$ $T_{3}$ $0.02196_{\pm 0.023}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0615}_{\pm 0.005}}0.0940\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}28.02_{\pm.78}}38.15\end{matrix}$ $98.90_{\pm.03}$ $86.1_{\pm.61}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}92.89}_{\pm.32}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 19\end{matrix}$
7 $\mathbb{R}^{3}$ none $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0219}_{\pm 0.0002}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}0.0672_{\pm}{0.001}}0.2051\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}27.54}_{\pm.68}}38.63\end{matrix}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}98.91}_{\pm.01}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.99}_{\pm.38}$ $90.05_{\pm.50}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 19\end{matrix}$
@@
EGNN - - - - 0.0290 0.0710 29.00 98.7 82.0 - -
DimeNet++ - - - - 0.0297 0.0435 24.60 - - - -
$\mathrm{SE}(3)$-T - - - - 0.1420 0.0510 53.00 - - - -
@@
Not a fair comparison as trained for longer. ∗∗Not a fair comparison, as MiDi uses open Babel optimization procedure.

### v2

@@ 4 Experiments
6 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SE}(3)$ $85.44_{\pm 0.03}$ $85.39_{\pm 0.11}$ $65.16_{\pm 0.47}$ $5.73_{\pm 0.14}$ $5.73_{\pm 0.13}$
7 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SE}(3)$ $84.75_{\pm 0.19}$ $84.65_{\pm 0.20}$ $4.77_{\pm 0.11}$ $4.78_{\pm 0.12}$
➟  8 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SE}(3)$ $85.48_{\pm 0.14}$ $85.41_{\pm 0.09}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.69}_{\pm 0.10}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.70}_{\pm 0.10}$
9 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $84.47_{\pm 0.15}$ $84.56_{\pm 0.02}$ $8.07_{\pm 0.30}$ $8.05_{\pm 0.28}$
10 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $85.38_{\pm 0.04}$ $85.36_{\pm 0.05}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}62.01}_{\pm 1.24}$ $5.36_{\pm 0.14}$ $5.37_{\pm 0.13}$
11 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $84.79_{\pm 0.16}$ $84.67_{\pm 0.21}$ $5.21_{\pm 0.18}$ $5.22_{\pm 0.17}$
12 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.76}_{\pm 0.04}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.74}_{\pm 0.05}$ $4.88_{\pm 0.04}$ $4.89_{\pm 0.04}$
@@
14 $\mathbb{R}^{3}\times S^{2}$ none $85.34_{\pm 0.17}$ $36.65_{\pm 2.52}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}61.87}_{\pm 0.36}$ $5.37_{\pm 0.06}$ $46.69_{\pm 0.27}$
15 $\mathbb{R}^{3}\times S^{2}$ none $84.93_{\pm 0.36}$ $33.35_{\pm 1.66}$ $5.80_{\pm 0.20}$ $58.66_{\pm 1.72}$
@@
16 $\mathbb{R}^{3}$ $T_{3}$ $85.20_{\pm 0.15}$ $31.58_{\pm 0.32}$ $65.81_{\pm 0.26}$ $5.56_{\pm 0.05}$ $52.68_{\pm 1.26}$
17 $\mathbb{R}^{3}$ $T_{3}$ $85.24_{\pm 0.07}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}33.11}_{\pm 0.81}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.21}_{\pm 0.18}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}39.15}_{\pm 0.69}$
18 $\mathbb{R}^{3}$ none $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.26}_{\pm 0.08}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}31.10}_{\pm 0.77}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}62.29}_{\pm 0.43}$ $6.24_{\pm 0.09}$ $77.98_{\pm 6.39}$
19 $\mathbb{R}^{3}$ none $84.29_{\pm 0.17}$ $32.97_{\pm 0.59}$ $5.99_{\pm 0.23}$ $>100$
Reference methods from literature (ShapeNet 3D)
@@
GeomGCNN - - - - - - - 89.10 - - -
Reference methods from literature (CMU Motion)
TFN - - - - - - - - - - 66.90 -
EGNN - - - - - - - - - - 31.70 -
CGENN - - - - - - - - - - 9.41 -
CSMPN ∗∗ - - - - - - - - - - 7.55 -
is trained with 1024 points, instead of 2048. ∗∗ uses additional pre-computed simplicial structures in training data.
4.1 3D point cloud segmentation & generation task
We evaluate our models on the ShapeNet 3D dataset (Chang et al., 2015) on part segmentation and generation tasks. We compare our models to various state-of-the-art methods like PointnXt (Qian et al., 2022), Deltaconv (Wiersma et al., 2022), and GeomGCNN (Srivastava & Sharma, 2021)for the part segmentation task. For the generation task, we compare to the latent diffusion model, LION, (Zeng et al., 2022), PVD (Zhou et al., 2021) and DPM (Luo & Hu, 2021). For additional details, see App. A.6.
Table 2: Ablation study on QM9 for property regression task using mean absolute error (MAE ×$\times 10^{-3}$) as the performance metric for three properties mu, alpha and $\epsilon_{HOMO}$, as well as molecule generation task (discovery) using atom stability, molecule stability and discover∗∗∗ as performance metric. Best results are highlighted in blue.
@@
Model Variation Type ! Coordinates as Scalars ! Coordinates as Vectors Effective Equivariance MAE $\mu$ (D) MAE $\alpha$ ($a_{o}^{3}$) MAE $\epsilon_{HOMO}$ (eV) Stab Atom % Stab Mol % Discover % Normalized Epoch Time
1 $\mathbb{R}^{3}$ $\mathrm{SE}(3)$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.41}_{\pm 0.37}$ $53.03_{\pm 0.92}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}22.29}_{\pm 0.27}$ $\sim 1$
2 $\mathbb{R}^{3}$ none $17.77_{\pm 0.21}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}52.22}_{\pm 0.93}$ $22.44_{\pm 0.15}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}96.74}_{\pm 0.07}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}72.39}_{\pm 0.50}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}89.06}_{\pm 0.39}$ $\sim 1$
@@
➟  4 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ $10.44_{\pm 0.22}$ $42.68_{\pm 1.04}$ $19.64_{\pm 0.75}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}99.38}_{\pm 0.02}$ $92.91_{\pm 0.41}$ $90.26_{\pm 0.20}$ $\sim 10$
5 $\mathbb{R}^{3}\times S^{2}$ none $10.53_{\pm 0.17}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}40.21}_{\pm 0.56}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}18.69}_{\pm 0.20}$ $99.33_{\pm 0.06}$ $92.71_{\pm 0.51}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}91.54}_{\pm 0.34}$ $\sim 10$
@@
6 $\mathbb{R}^{3}$ $T_{3}$ $22.11_{\pm 1.03}$ $61.97_{\pm 1.05}$ $25.83_{\pm 0.25}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}98.57}_{\pm 0.01}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}81.62}_{\pm 0.15}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}91.83}_{\pm 0.45}$ $\sim 1$
7 $\mathbb{R}^{3}$ none $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.40}_{\pm 0.37}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}58.18}_{\pm 0.09}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}24.54}_{\pm 0.26}$ $98.40_{\pm 0.02}$ $78.48_{\pm 0.36}$ $89.26_{\pm 0.30}$ $\sim 1$
@@
EGNN - - - - 29.0 71.0 29.0 98.7 82.0 - -
DimeNet++ - - - - 29.7 43.5 24.6 - - - -
$\mathrm{SE}(3)$-T - - - - 53.0 51.0 53.0 - - - -
@@
Not a fair comparison as trained for longer. ∗∗Not a fair comparison, as MiDi uses open Babel optimization procedure. ∗∗∗It is defined as a product of validity $\times$ unique $\times$ novelty.